OpenStudy (anonymous):

Check my answer plz Short Answer: Raise the quantity in parentheses to the indicated exponent, and simplify the resulting expression. Express answers with positive exponents. (-36x^2y/180x^-3y^-5)^-2

OpenStudy (anonymous):

Following the laws of exponents... (ab)ⁿ = aⁿbⁿ a^n/a^b = a^(n - b) We obtain: (-36x²y/180x^-3y^-5)² => (-36x²y)²/(180x^(-3)y^(-5))² => 1296x^4y²/(32400x^(-6)y^(-10)) => x^4y²/25x^(-6)y^(-10) => x^(4 - (-6))y^(2 - (-10))/25 => x^(10)y^(12)/25

OpenStudy (angela210793):

isn't the exponent -2?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

you have it upside down. everything is correct except you ignored the minus sign

OpenStudy (anonymous):

should be \[\frac{25}{x^{10}y^{10}}\]